A 13.0 cm -diameter wire coil is initially oriented so that its plane is perpendicular to a magnetic field of 0.51 T pointing up. During the course of 0.16 s , the field is changed to one of 0.20 T pointing down. What is the magnitude of the average induced emf in the coil?
Answer #1

We know; induced emf is given by

e=d/dt

Now here = B. A

Where B= magnetic field and A= area of the coil

So, e=A*(dB/dt) ; where A= 3.14×(6.5×10-2 )2

=3.14×(6.5×10-2 )2 ×(.51-(-.20))/(.16) V/m

=5.89×10-4 V/m